Assume no taxes. Storm Inc. has 250,000 shares outstanding, P=$20, no debt. Current WACC is 15%. As of now, management expects that EBIT will be $750,000 per year forever. It now plans to buy another firm ABC Inc. (in the same industry, so assume similar risk) at a cost of $300,000. ABC will add $120,000 per year in perpetuity.

a) Suppose that Storm issues equity to buy the other firm. What will happen to Storm’s firm value when the merger is announced? Equity value? Share price?

b) How many shares will Storm have to issue to fund this purchase?

c) What is the return to Storm’s equityholders or cost of unlevered equity? Confirm your answer using the PV of all cash flows to shareholders at unlevered equity cost.

Repeat the analysis assuming that Storm issues debt (cost of 10%) to buy the other firm.

a) What will happen to Storm’s firm value when the merger is announced? Equity value? Debt value? Share price?

b) How much debt will it issue to fund the purchase?

c) What is the cost of levered equity? Confirm your answer by both computing the PV of all cash flows to shareholders at levered equity cost and using MMII.
Answer #1

Ans a) If Storm issues to buy the other firm

1 ) Value of firm= EBIT/ WACC

=750000+120000/0.51= 5800000

EQUITY Value= Value of firm- value of debt

= 5800000-0=5800000

Calculation of share price = Equity value/ total Amountof shares Outstanding

= 5800000/250000+300000

=$10.54

B) no of shares to be issued

Total cost of purchase of new firm/ share price

=300000/10.54

=28463 shares

C) calculation of return to equityshare holders

= Net income available for equityshare holders/ no of shares outstanding

No of shares outstanding=12500 shares of $20 each +28463 shares of RS. 10.54 each

= 870000/40963

=$21.24 per share

B) if Storm issues debt for purchasing a new firm

AMOUNT
EBIT 870000
Less : interest on debentures@ 10% 30000
Earning available for Equity shareholders 840000
Total cost of capital (given) 15%
Value of the firm = EBIT/Cost of capital= 870000/.15 5800000

C) calculation of cost of equity

Ke= Earnings available for Equity shareholders/value of equity

Amount
Value of debt 300000
Value of Equity= value of firm- value of debt= 5800000-300000 5500000

Ke= 840000/5500000=15.27%

Calculation of share price = value of equity/ amount outstanding

= 5500000/250000

$22 per share

Calculation of ke by mmii approach

Value of equity = Net operating income- interest/cost of equity

5500000=840000/cost of equity

Cost of equity= 840000/5500000

=15.27%